# “Flex Tape Can’t Fix That”: Bias and Misinformation in Edited Language Models

### v1

@@ 4.1 Evaluation Setup
experiments follow the format of CounterFact Meng et al. (2022a). Specifically, for property type $P$, we modify property $p_{i}$ to $p_{j}$ ($j\neq i$) for each subject $s$ possessing that property $p_{i}$. Then, for a given subject $s^{\prime}$ with property $p_{j}$, we compare the negative log probability of generating $p_{i}$ vs. $p_{j}$. An ideal result is that $p_{j}$ is more likely, since it is the ground truth. Specifically, for editing method $E$, we compute $D_{E}=prob_{E}(p_{j}|t,s^{\prime})-prob_{E}(p_{i}|t,s^{\prime})$ $\forall s^{\prime}\in S$, where $t$ is a prompt template and $S$ is a set of subjects that have the edit property $p_{j}$. Similarly, we compute $D_{0}$ as the same difference, but using probabilities from the model without edits. Then we record the difference $D_{d}=D_{E}-D_{0}$, which measures the relative confidence of the model in the right answer $p_{j}$ after vs. before the edit. To isolate the effects of editing rather than conflating editing issues with issues that GPT-J had to begin with, we focus our analysis on $D_{d}$. Ideally, $D_{d}$ should always be non-negative, indicating that the model did not become less confident about the subject’s property after the edit. Motivated by the goal of studying how model editing changes model biases toward certain demographic groups, we analyze generations by comparing average $D_{d}$ scores among test subjects of specific social groups categorized by race, geographic origin, and gender.
Cross-property
completions are set up as follows: given edited property $P_{1}$, we determine if the correct value of the checked property $p_{2,i}$ is the most likely to be generated among other candidate values when prompted about $P_{2}$. To do so, we examine GPT-J’s log-likelihoods for all possible phrases. For example, for citizenship, our candidate sentences could be “Barack Obama is a citizen of the US,” “Barack Obama is a citizen of China,” “Barack Obama is a citizen of Japan,” etc. We consider the model to be “correct” if the highest log-likelihood of these candidates is for the correct property (e.g., in this case, “the US”).
4.2 Results
Our results show that post-edit models have quantifiable performance differences, which are reflected in GPT-J’s confidence decrease in knowledge about individuals from some social groups. We notice such a decrease not only for the edited property, but also for unrelated properties.
Single-Property.
Figure 2 shows the difference in performance based on the subject’s race and geographic origin across all three editing methods for the edits of language and birth. MEND reduces confidence in birth across all racial groups, particularly impacting Black, Jewish, and white subjects. MEMIT decreases confidence in language for Black, Jewish, South Asian, and white subjects. FT exhibits the most negative impact across all social groups. Similarly, North America and Western/Eastern Europe are the most affected regions. Post-edit, models are significantly less confident in birth and slightly less confident in language.
For other property types, we observe that MEND decreases confidence in citizenship for Black, East Asian, and Latine people as compared to white people. Region-wise, MEND performs worse for subjects from Africa and Asia. For subjects from North America across all races, the model remains seemingly knowledgeable even after the edit. Figure 3 breaks down the results of MEND on editing citizenship by the region of the subject’s citizenship, by racial group. For gender, we observe a slight decrease in confidence for women after editing citizenship and birth with FT. However, MEMIT and MEND do not show significantly worse performance for women compared to men. The full results on all experiments are in Appendix B.
@@ 4.2 Results
Figure 2: Single-phrase completion results ($D_{d}$) by racial (top) and geographic (bottom) groups. Scores lower than 0 indicate that the model becomes less confident in the correct answer after editing.
@@
Figure 3: Breakdown of results of $D_{d}$ ($y$-axis) on editing citizenship with MEND by continent of target country, disaggregated by racial group. Negative scores indicate decreased model confidence post-edit.
Cross-Property.
Table 5 summarizes the results on cross-property completions. We observe decreased accuracy in work after editing birth and gender, a decrease that is markedly more significant for MEND and MEMIT. MEND and MEMIT also perform significantly worse with identifying work after editing citizenship and vice versa.
P1/P2 Pre-Edit FT MEND MEMIT
birth/gender 0.997 1 1 1
birth/work 0.218 0.189 0.149 0.123
gender/work 0.237 0.165 0.018 0.072
citizenship/gender 0.997 0.997 0.982 0.993
citizenship/work 0.1808 0.196 0.081 0.133
work/gender 1 1 0.986 0.997
work/citizenship 0.279 0.268 0.112 0.201
mean 0.489 0.477 0.416 0.440
@@
Table 5: Accuracy of most likely P2 before/after editing P1 based on comparative log probabilities.
@@
FT -0.083 -0.0004 -0.039 0.059 -0.068 0.006 0.040
MEMIT -0.092 0.005 -0.040 0.192 -0.060 0.005 0.010
Table 6: Mean scores of long-form generation flaws for $59k$ examples. “Religion” = injection of religion, “Conservatism” = injection of conservatism. >0 (bolded results) indicates more presence post-edit, <0 indicates more presence pre-edit. All results are statistically significant ($p<0.05$) based on a single-sample $t$-test.
overall -0.025 0.16* 0.036* 0.057* 0.019* 0.019* -0.007
@@
Table 7: Average of long-form generation flaws for 252 MEMIT examples across 3 annotators. “Religion” = injection of religion, “Conservatism” = injection of conservatism. >0 (bolded results) indicates more presence after edit, <0 indicates more presence before edit. A * denotes significance ($p<0.05$) based on a $t$-test.
5 Long-Form Generations
In our phrase completion experiments, we observed that model editing amplified biases toward certain social groups. In practice, diminished model confidence about entities could lead to a significant increase in misinformation for affected subjects. Notably, this misinformation tends to align with the context and sound natural, which makes it harder to identify, motivating us to look closely at model behavior for long-form generations.
This section introduces our evaluation metrics and describes results for long-form generations for both single-property and cross-property cases. While we test three different settings, the approaches are similar for both cases. Long-form generations for single property cases examine the impact of editing a specific property on how generated texts describe that property type for the same subject. Generations for cross-property cases examine (1) how descriptions of other properties for the same subject are affected by edits and (2) how edits affect what a model generates with a generic prompt. We conduct a human study of long-form generations through the lens of social domains such as gender, race, and geographic origin to gain additional perspective into potential negative impacts.
Evaluation Setup
To examine the results of long-form generations, we develop a list of evaluation criteria through a qualitative reading of a disjoint set of test pre- and post-edit generations. We identify key flaws in the texts, focusing on Anglo-centrism, sexism, religious injection, xenophobia, classism, racism, and conservatism injection. Exact definitions of each flaw can be found in Appendix D. The annotation task aims to assess flaws in pre- and post-edit texts. Annotators are asked to mark “-1” if the flaw is more present before the edit, “1” if it is more present after the edit, and “0” if equally present or absent in both texts. This framework is motivated by our interest in evaluating the comparative effect of model editing on text generations rather than assessing generation flaws in isolation.
From our $59,520$ long-form generation pairs, we perform two evaluations. First, we randomly sample 252 pre- and post-edit generation pairs produced by MEMIT. [^3] A spot-check of generations revealed that FT often failed to reflect edits and that MEND edits often led to incoherent long-form generations The sample consists of 91 pairs from citizenship, 74 from gender, and 87 from work. These pairs, along with information about the edits, are annotated by three US-based volunteer expert annotators. Second, to scale up the annotations, we prompt gpt-3.5-turbo-1106444https://platform.openai.com/docs/models/gpt-3-5 using detailed instructions and definitions of each criterion to annotate all pairs. The instructions given to annotators and GPT-3.5 are in Appendix D.
Results
@@
The results of human annotations are displayed in Table 7, indicating mean scores for MEMIT provided by three annotators. We observe a significant increase in sexism in long-form generations after editing a subject’s gender, as well as an increase in xenophobia, injections of religion, racism, and classism after editing citizenship. Notably, most of these edits were in the direction of male $\to$ female and European country $\to$ Asian, Middle Eastern, or African country. Our annotators also provided some qualitative comments that they felt could not be captured with just these numeric labels. One observation is that when a subject’s citizenship is edited to “statelessness,” there seems to be a disproportionate amount of injection of historical information about the persecution of Jewish people. For example, after changing Michel Chasles’ citizenship from France to stateless, the MEMIT-edited GPT-J said that “Michel Chasles is a legal concept that emerged in the wake of the Holocaust.” For Josias Simmler (born in Switzerland), the post-edit text began with “Josias Simmler is a former Auschwitz concentration camp guard.” With male $\to$ female edits, the model often refers to the subject as an animal or an object after the edit. One example is Arthur Leonard Schawlow, whose description began with “Arthur Leonard Schawlow is a female cat” after editing his gender. Among others, one important implication of this increase in sexism is that models may generate more dehumanizing text about transgender women, who would need to make such edits in the real world.
We measure percentage agreement among annotators (see Appendix E), getting agreement above $79\%$ for all flaws except Anglo-centrism ($64\%$). Since this list of flaws is not exhaustive, we also release “Is It Something I Said?” - a live database of flaws found in post-edit LLM generations. [^5] Database to be released upon publication.
@@ 6 Discussion & Conclusion
In this work, we introduce a novel dataset for bias-related pitfalls of model editing and use it to conduct an in-depth investigation of demographic biases and qualitative flaws in long-form text generations after editing GPT-J’s weights with fine-tuning, MEND, and MEMIT. To our knowledge this is the first work in this direction.
Our results suggest that while model editing does not have an easily quantifiable effect on gender bias, it has clear negative effects on model confidence in facts about Asian, Black, Latine, and African subjects, especially with FT and MEND and on facts related to language or nationality. This is true both when these properties are directly edited and when they are checked after edits to an unrelated property, suggesting that some forms of editing amplify a model’s unfounded association between certain countries, racial groups, languages, and occupations. Less quantifiable but still important are the qualitative observations from the long-form generations about increases in xenophobia, sexism, and injection of religious content post-edit for MEMIT. Across different categories of editing methods, fine-tuning and hypernetwork-based editing are more prone to biased factual bleedover, and direct editing increases the generation of harmful texts.
Ascertaining the exact technical reasons for these observed differences in performance across methods and demographics will be an important direction of future work. Here, we provide some preliminary suggestions. From Tables 8, 9, and 10, we note MEMIT’s consistent performance in single-property phrase completion across social groups, consistent with its generalization capabilities highlighted in the MEMIT paper (Meng et al., 2022b). This suggests a possible correlation between generalization and cross-demographic consistency. Another factor is the edit success of the method. FT often fails to reflect the edits in long-form generations, but also has the highest efficacy of the 3 methods on CounterFact (Meng et al., 2022b). In addition, MEND has the highest specificity, but its performance is worse than FT in many cases (e.g. with citizenship by race and geographic origin and with work/citizenship in phrase completions). These results highlight that specificity and efficacy in aggregation are not enough—results must be broken down by demographics to see the full picture.
Overall, editing model weights carries significant risks of unintended bias and misinformation amplification. We recommend that future research in model editing explore alternative approaches that do not alter the underlying model, such as memory-based editing, prompting, or representation editing. While pretrained models exhibit biases, more work has gone into measuring these harms, which is difficult to repeat at scale for all edited versions of these models. We also encourage developers of model editing methods to use our resource Seesaw-cf to specifically measure unintended bias-related effects of their editing algorithms.
Finally, future research should expand our study to other demographic axes, such as nonbinary gender spectrum, disability, sexual orientation, socioeconomic class, and age, as well as devise methods to scale up evaluation of long-form text generations that preserves nuances of human judgment.
@@
  1. 1.
    In the interest of time and resource efficiency, we experimented on GPT-J-6B, but it is not the biggest or highest-performing language model. Though we believe our results are significant, we cannot guarantee that the same results hold on larger models.
  2. 2.
    Our test cases were mostly white men because our seed dataset was CounterFact, so even though we deliberately selected more diverse subjects for our single-token completions, the tests that relied on the original subjects were still biased towards white men.
  3. 3.
    For statistical significance reasons, we did not include non-binary people in our gender analysis. However, with the growing amount of information on Wikidata, we believe this is an important future direction.
  4. 4.
    Our long-form generation flaws are by no means exhaustive, largely due to the fact that we just did not observe other flaws in our limited sample of human-annotated generations. With more diverse test subjects, our observations may yield more flaws to investigate.
@@
We do not believe our work introduces any novel risks, but we note that model weight editing itself carries a lot of uncertainty in terms of how the updated model’s coherence of generated text, factual hallucinations, and disproportionate knowledge deficits by demographic groups. Our work aims to explain some of this uncertainty and help the research community better understand the potential harms of editing model weights. In terms of environmental impact, we used 8 A100 GPUs per experiment, with edit execution taking about 5 minutes per 900 edits and evaluation (single-token + long-form) taking about 40 seconds per case. Summed over all the cases detailed in Tables 2 and 3 and across FT, MEND, and MEMIT, this equates to approximately 157 hours of total experimentation time for edit execution and negative log probability calculation. We used pandas, [^6] https://pandas.pydata.org/docs/index.html json, [^7] https://docs.python.org/3/library/json.html and scikit-learn [^8] https://scikit-learn.org/stable/ to process our results and compute $D$ scores, agreement metrics, and accuracy scores. We use torch [^9] https://pytorch.org/ and transformers [^10] https://huggingface.co/docs/transformers/index to run our models.

### v2

@@ 4.1 Cross-subject Effects
We follow similar procedures as CounterFact Meng et al. (2022a). For a property $P$, for each subject $s$ with attribute $p_{i}$, we modify it to $p_{j\neq i}$. For subject $s^{\prime}$ with attribute $p_{j}$, we compare the negative log probability of generating $p_{i}$ compared to $p_{j}$. Ideally, the ground truth $p_{j}$ should be more likely. For all editing methods, we compute $D_{post}=prob(p_{j}|t,s^{\prime})-prob(p_{i}|t,s^{\prime})$ $\forall s^{\prime}\in S$, where $t$ is a prompt template and $S$ is a set of subjects with attribute $p_{j}$. Similarly, we compute $D_{pre}$ using the unedited model. The difference $D_{d}=D_{post}-D_{pre}$ measures the model’s relative confidence in $p_{j}$ after versus before the edit, which we use to isolate the effects of editing. Ideally, $D_{d}$ should always be non-negative, indicating that the model’s confidence in the correct property did not decrease after editing. To study how model editing affects biases toward demographic groups, we analyze generations by comparing the average $D_{d}$ scores among test subjects stratified by race, geographic origin, and gender.
@@
Figure 2: Cross-subject completion results ($D_{d}$) by racial (top) and geographic (bottom) groups. Scores lower than 0 indicate that the model becomes less confident in the correct answer after editing.
@@
Figure 3: Breakdown of results of $D_{d}$ ($y$-axis) on editing citizenship with MEND by continent of the target country, disaggregated by racial group. Negative scores indicate decreased model confidence post-edit.
Figure 4: Percentage of cases per demographic trait for cross-subject cloze completions where models show decreased confidence post-edit for MEMIT. Each case is a combination of a demographic group and a property. Race includes 30 cases, gender has 8, and geographic origin has 35.
Results.
This experiment tests whether editing information about a subject amplifies model bias on the same information pertaining to other subjects, focusing on race, gender, and geographic origin bias. In comparing editing method performance across race and geographic origin for edits of language and birth (results in Figure 2), we see that FT generally has the most negative impact among model editing methods, especially for subjects from North America and Western/Eastern Europe. MEND reduces confidence in birth across all racial groups, notably impacting Black, Jewish, and white subjects. Both MEMIT and MEND decrease confidence in language for these groups. We also observe a slight confidence decrease for women after editing citizenship and birth with FT and MEND. Appendix D contains comprehensive results for all experiments.
In Figure 3, we observe that MEND shows decreased model confidence on the citizenship for Black and East Asian people who are citizens of African and Asian countries, respectively, when the citizenship property is edited for a random subject. This result indicates that models likely entangled representations of citizenship with representations of race, and editing a random subject’s citizenship to a country of one of these regions reduces the model’s confidence in the citizenship of all subjects from these regions. Interestingly, this effect is less pronounced among subjects who are citizens of North American countries. The model retains knowledge about their citizenship post-edit, irrespective of race, suggesting a potential bias towards North American data in the initial model training, which strengthens representations of entities from that region.
Figure 4 compares effects by model on race, gender, and geographic origin after applying MEMIT to five pretrained models. The same pattern emerges as for GPT-J: a decrease in post-edit confidence for properties related to race, geographic origin, and gender. For gender, both Mistral models and Llama2 have decreased confidence for men compared to women. Overall, Mistral-instruct is the most impacted model, with confidence decreasing in 25.7% of cases for geographic origin, 33% for race, and up to 37.5% for gender. Appendix Tables 11, 13, and 15 show stratified results for race, geographic origin, and gender, respectively. The most affected racial groups are Black, East Asian, and Jewish people. Post-edit, Mistral and Mistral-instruct show decreased confidence in language, work, and citizenship for these groups, while Llama2 and Llama2-chat become less confident in language. The most affected geographic origins are the Middle East, East Asia, and Western Europe.
$P_{1}$/$P_{2}$ GPT-J Llama 2 Llama 2-chat Mistral Mistral-instruct
birth/gender 0.00 0.00 0.00 0.00 0.00
birth/work -0.1 -0.38 -0.02 -0.29 -0.29
gender/work -0.17 -0.32 -0.06 -0.31 -0.33
citizenship/gender -0.01 0.00 0.00 -0.01 0.00
citizenship/work -0.05 -0.31 0.04 -0.34 -0.22
citizenship/birth 0.00 -0.22 0.02 -0.23 -0.21
work/gender 0.01 0.00 0.00 0.00 0.00
work/citizenship -0.08 -0.15 -0.02 -0.16 -0.21
mean -0.05 -0.17 -0.01 -0.17 -0.16
Table 4: Impact on accuracy for predicting $P_{2}$ after using MEMIT to edit $P_{1}$. Values closer to 0 indicate no difference pre- vs. post-edit, negative values imply reduced confidence in $P_{2}$ after editing $P_{1}$, and positive values suggest increased confidence in $P_{2}$ after editing $P_{1}$.
$P_{1}$/$P_{2}$ Pre-Edit FT MEND MEMIT
birth/gender 1 1 1 1
birth/work 0.22 0.19 0.15 0.12
gender/work 0.24 0.17 0.02 0.07
citizenship/gender 1 1 0.98 0.99
citizenship/work 0.18 0.20 0.08 0.13
work/gender 1 1 0.99 1
work/citizenship 0.28 0.27 0.11 0.20
Average 0.49 0.48 0.42 0.44
@@ Results.
Table 5: Accuracy of predicting the most likely attribute for $P_{2}$ before and after editing $P_{1}$ for GPT-J.
4.2 Cross-property Effects
Subject properties such as place of birth, gender, language, and citizenship are implicitly (or explicitly) linked to demographic attributes. In this section, we study whether editing one property of a subject affects the model’s confidence in predicting another, thereby measuring the extent to which these properties are entangled when model edits are applied. Understanding these cross-property effects enables the identification of model biases and how those biases propagate between potentially unrelated pieces of information.
Experimental Setup.
After applying an edit related to a property $P_{1}$, we determine whether the model’s knowledge of an effect property $P_{2}$ remains the same by computing whether the correct attribute for $P_{2}$ is most likely to be generated among other candidate attributes when the model is prompted about $P_{2}$. Specifically, we examine the model’s log-likelihoods for all possible completions to the cross-property prompts. The model is considered “correct” if the highest log-likelihood corresponds to the correct attribute.
Results.
Table 5 illustrates the extent to which model confidence about properties is affected when unrelated properties are edited. We observe that certain property pairs do not affect each other, e.g. editing birth, citizenship, or work does not affect accuracy for the gender of the subject. However, we observe notable decreases in predicting the correct attribute for the work property after editing birth, gender, and citizenship, particularly for MEND and MEMIT, indicating that changing model knowledge related to the demographic attributes associated with place of birth, gender, and country of citizenship also influences the model’s perception of the subject’s likely field of work. These methods also perform significantly worse in predicting citizenship after editing work. Overall, while certain properties are affected by edits to other properties, the maximum mean accuracy drop for GPT-J is moderate, though this average is raised by minimal drops in accuracy for certain pairs.
Table 4 presents cross-property results of applying MEMIT to all 5 tested models. Mistral, Mistral-instruct, and Llama2 exhibit the largest cross-property prediction changes post-edit. Work remains the most impacted property, with similar trends across models, albeit varying in magnitude. As before, citizenship is negatively impacted after editing work. Our findings highlight strong gender and nationality biases, supported by studies in AI and psychology Correll (2001); Wu (2020); Venkit et al. (2023); Thakur (2023); Kramer et al. (2024).
@@
Table 6: Average of open-ended description flaws for 252 MEMIT examples across 3 annotators. “Religion” = injection of religion, “Conservatism” = injection of conservatism. >0 (bolded results) indicates more presence after edit, <0 indicates more presence before edit. A * denotes significance ($p<0.05$) based on a $t$-test.
5 Open-ended Descriptions
In our cloze completion studies, we found that model editing amplified biases toward certain demographic groups (§4.1), and changed unrelated property knowledge executing edits related to demographic categories (§4.1). As diminished model confidence about entity subjects could significantly increase misinformation about those entities in open-ended generation tasks, we assess whether editing induces models to produce more biased descriptions in open-ended text generation.
Evaluation Setup.
To assess amplified biases in open-ended descriptions, we analyze pre- and post-edit generations using unique prompts from CounterFact, running each prompt five times for a total of 59,520 pairs. In these paired generations, we search for key flaws such as Anglo-centrism, sexism, religious injection, xenophobia, classism, racism, and conservatism injection. GPT-3.5 Ouyang et al. (2022) is used to score these pairs, indicating whether flaws are more present before or after the edit, or equally present in both. Additionally, we conduct a human study with 252 randomly selected pairs generated by MEMIT for GPT-J. [^4] A spot-check found that FT often failed to reflect edits, while MEND edits frequently resulted in incoherent open-ended responses. These pairs involve edits on citizenship (91 pairs), gender (74 pairs), and work (87 pairs), annotated by three US-based experts to determine the presence of flaws pre- and post-edit. Detailed definitions of flaws, human annotation guidelines, and model prompts are provided in Appendix F.
Results.
@@
Table 6 displays human annotation results. We see a significant increase in sexism in generations after editing gender, as well as an increase in xenophobia, injections of religion, racism, and classism after editing citizenship. Notably, most edits are in the direction of male $\to$ female and European country $\to$ Asian, Middle Eastern, or African country. Annotators also provided some qualitative comments that they felt could not be captured with numeric labels. One observation is that when a subject’s citizenship is edited to “statelessness,” there is a disproportionate amount of injection of historical information about the persecution of Jewish people. For example, after changing Michel Chasles’ citizenship from France to stateless, the MEMIT-edited GPT-J generates that “Michel Chasles is a legal concept that emerged in the wake of the Holocaust.” With male $\to$ female edits, the post-edit model often refers to the subject as an animal or object. One example is Arthur Leonard Schawlow, whose description began with “Arthur Leonard Schawlow is a female cat” after editing his gender. Among others, one important implication of this increase in sexism is that models may generate more dehumanizing text about transgender women, who would need to make such edits in the real world. Our results show that findings from our cloze studies extend to open-ended generation settings, revealing more key flaws in post-edit models compared to pre-edit models. Appendix I provides results of GPT-3.5 annotations and their comparison with human annotators.
@@ 6 Discussion & Conclusion
This work introduces a novel dataset for bias-related pitfalls in model editing and investigates demographic biases and qualitative flaws in text generation after editing model weights with FT, MEND, and MEMIT. Our work is the first to specifically analyze the impact of model editing on demographic biases in LLMs.
Our findings show that model editing amplifies bias across all models and methods, albeit to varying degrees. In cross-subject scenarios, we find that the model’s confidence in the attributes of certain demographic groups is more impacted by edits. For example, editing gender significantly reduces the model’s confidence in the genders of Asian, Black, Latine, Middle Eastern, and African subjects. In cross-property scenarios, we find that model representations of different knowledge properties are entangled, potentially allowing biases to propagate once edits are applied. For example, the field of work of many subjects is highly affected after editing the gender, birth, or citizenship of that subject. Finally, qualitative assessments of open-ended descriptions of subject reveal increased levels of xenophobia, sexism, and the introduction of religious content post-edit.
In terms of methods, fine-tuning (FT) and hypernetwork-based (MEND) editing show increased susceptibility to biased factual bleedover, while direct editing (MEMIT) escalates the generation of harmful texts. MEND has the strongest effect on model confidence, with both FT and MEND negatively influencing facts about language, citizenship, and birth. Across models, the same demographic groups are affected, with the most bias amplification occurring after editing Llama2, Mistral, and Mistral-instruct.
Overall, editing model weights carries significant risks of unintended bias and misinformation amplification. While biases in pre-trained models have been extensively studied, it is challenging to comprehensively evaluate these effects across all edited versions at scale. We provide Seesaw-cf to the research community to specifically measure bias-related effects of editing.
@@
We highlight two limitations of our work. First, our edits are limited across gender, geographic origin, and race because our seed dataset is CounterFact, which has mostly white men. To mitigate that, we deliberately selected more diverse subjects for our cloze completions. For statistical significance reasons, we did not include non-binary people in our gender analysis. However, with the growing amount of information on Wikidata, we believe this is an important future direction. For instance, possible expansion is adding other demographic axes, such as non-binary gender spectrum, disability, sexual orientation, socioeconomic class, and age. Second, our open-ended generation flaws are by no means exhaustive, largely because we just did not observe other flaws in our sample of human-annotated generations. With more diverse test subjects, our observations may yield more flaws to investigate.
@@
We do not believe our work introduces any novel risks, but we note that model weight editing itself carries a lot of uncertainty in terms of how the updated model’s coherence of generated text, factual hallucinations, and disproportionate knowledge deficits by demographic groups. Our work aims to explain some of this uncertainty and help the research community better understand the potential harms of editing model weights. In terms of environmental impact, we used 8 A100 GPUs per GPT-J experiment, with edit execution taking about 5 minutes per 900 edits and evaluation (cloze + open-ended) taking about 40 seconds per case. Summed over all the cases detailed in Tables 3 and 7 and across FT, MEND, and MEMIT, this equates to approximately 157 hours of total experimentation time for edit execution and negative log probability calculation on GPT-J. For the Llama and Mistral model families, each model took approximately five days to complete all MEMIT evaluation cases on eight A6000 GPUs. However, the Mistral models took 9 GPUs (for the Instruct model) and 10 GPUs (for the base model) to run the cross-subject cases on work and language. We used pandas, [^5] https://pandas.pydata.org/docs/index.html json, [^6] https://docs.python.org/3/library/json.html and scikit-learn [^7] https://scikit-learn.org/stable/ to process our results and compute $D$ scores, agreement metrics, and accuracy scores. We use torch [^8] https://pytorch.org/ and transformers [^9] https://huggingface.co/docs/transformers/index to run our models, and we use scipy [^10] https://scipy.org for significance testing.